OpenStudy (anonymous):

the sum of 5 and a number is 6 x+5/x=6 x^2-6x+5=0 (x-5)(x-1)=0 x=5 or 1 how do i check this

OpenStudy (lgbasallote):

you check it by plugging in the value of x you got into every x in the equation...if the left hand side equals the right then it's right

OpenStudy (lgbasallote):

is the question: the sum of a number and the quotient of 5 and that number is 6?

OpenStudy (lgbasallote):

otherwise your interpretation is wrong

OpenStudy (anonymous):

no it the sum of 5 and a number is 6

OpenStudy (lgbasallote):

the sum of 5 and a number is 6 means x+ 5 = 6

OpenStudy (anonymous):

yes i got that but iam suppose to solve it then check my answer

OpenStudy (lgbasallote):

x + 5 = 6 x = 6 - 5 x = 1 \(\checkmark\)ing: 1 + 5 = 6? 6 = 6 \(\checkmark\)

OpenStudy (anonymous):

sweet i still feel lost i kinda had youo at putting 5 in where the xs are and it equaled out right